Using the following equations for force, work done, power, current and potential difference (or otherwise) give the following derived units in terms of base units. Show all of your working.

F=ma, W=Fd, P=ΔW/Δt', I=ΔQ/tΔ, V=W/Q

a. Joules

b. Watts

c. Volts


1
Expert's answer
2022-05-30T15:18:22-0400

(a)

"\rm J=N*m=kg*m/s^2*m=kg*m^2/s^2"

(b)

"\rm W= J/s=kg*m^2/s^3"

(c)

"\rm V=\frac{J}{C}=\frac{kg*m^2/s^2}{A*s}=\frac{kg*m^2}{A*s^3}"